OpenStudy (spring98):

What is the distance between A and B? Round your answer to the nearest hundredth. A. 9.58 B. 10.63 C. 11.56 D. 12.29

OpenStudy (jdoe0001):

well, if we knew either, that'd help otherwise I'd have to rub my crystal ball but I think that'd just make it shinier only

OpenStudy (anonymous):

\[\sqrt{(x2-x1)^2 + (y2-y1)^2}\]

OpenStudy (jdoe0001):

well, what are coordinates for point A and B?

OpenStudy (spring98):

(0, 6) and (7,-2)

OpenStudy (jdoe0001):

well then \(\bf \textit{distance between 2 points}\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ % (a,b) A&({\color{red}{ 0}}\quad ,&{\color{blue}{ 6}})\quad % (c,d) B&({\color{red}{ 7}}\quad ,&{\color{blue}{ -2}}) \end{array}\qquad % distance value d = \sqrt{({\color{red}{ x_2}}-{\color{red}{ x_1}})^2 + ({\color{blue}{ y_2}}-{\color{blue}{ y_1}})^2}\)

OpenStudy (anonymous):

0 = x1 7 = x2 6 = y1 -2 = y2 Plug into formula above

OpenStudy (anonymous):

\[\sqrt{(0-7)^2+(6-(-2))^2}\]

OpenStudy (spring98):

umm so would the answer be B?

OpenStudy (anonymous):

yes

OpenStudy (spring98):

@jdoe0001 is @Lily2913 right?

OpenStudy (spring98):

when someone helps me i like my answers to be confirmed.

OpenStudy (anonymous):

Yes B is right

OpenStudy (jdoe0001):

well, what did YOU get though?

OpenStudy (spring98):

thanks you! I wish i could medal you all for helping. :)